the cross-sectional area of an adjustable single loop is reduced from 1.0m^2 to 0.50m^2 in 0.10s. What is the average emf that is induced in this coil if it is in a region where B=2.0 T upward, and the coil's plane is perpendicular to B?

1
Expert's answer
2022-05-16T13:40:54-0400
"V=-\frac{\Delta \Phi}{\Delta t}=-\frac{B\Delta A}{\Delta t}"

"V=-\frac{2.0*(0.5-1.0)}{0.10}=10\:\rm V"